Assume the annual salaries of police officers are normally distributed with a mean of ​$47 000 and a standard deviation of ​$8000. Determine the percent of police officers with an annual salary of at least ​$47,000 Click here to view page 1 of the standard normal distribution table.... Click here to view page 2 of the standard normal distribution table.... Determine the percent of police officers with an annual salary of at least ​$47,000 . ​(Round to two decimal places as​ needed
Answer #1

Solution :

Given that ,

mean = = $47000

standard deviation = = $8000

P(X >$470000 ) = 1 - P(x < 47000)

= 1 - P((x - ) / < ( 47000 - 47000) /8000 )

= 1 - P(z <00 )   Using standard normal table,

= 1 - 0.5   

= 0.5

answer = 50.00%.